OpenStudy (anonymous):

Find the domain, intercepts, the intervals where function is increasing and decreasing, extreme values, intervals of concavity, points of inflection, asymptotes (vertical, horizontal, or slant). Show the1st and the 2nd derivative tables. Sketch a graph. f(x)= x^2+12/(x-1)

OpenStudy (anonymous):

sorry it wouldnt let me make an equation but its (x^2+12)/(x-1)

OpenStudy (anonymous):

\[\huge f(x)=\frac{x^2+12}{x-1} \]

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

find f'... set =0, solve for critical numbers....

OpenStudy (anonymous):

(x^2-2x+12)/(x-1)^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\large f(x)=\frac{x^2+12}{x-1} \] \[\large f'(x)=\frac{x^2-2x-12}{(x-1)^2} \]

OpenStudy (anonymous):

sorry jeannie... gotta go.... bye...

OpenStudy (anonymous):

ahhh!

OpenStudy (paxpolaris):

\[\large f(x) = {x^2+12 \over x-1}= {(x^2-1)+13 \over x-1}=\Large (x+1)+{13\over x-1}\] so the asymptotes are “\(y=x+1\)” and “\(x=1\)” you need the second derivative to show when the graph is concave up or down... although you can guess easily.

OpenStudy (anonymous):

thanks sorry i had to go get my car inspected